OpenStudy (yuii):

What linear function can be represented by the set of ordered pairs? {(−1, −10),  (3, 2),  (5, 8),  (7, 14)} f(x) =

OpenStudy (karim728):

first we know its a linear function.. therefore y=mx +b. take any two points and find the slop, then use y-y1=m(x-x1)